OpenStudy (anonymous):

evaluate the dot product of (8,4) and (-3,5)

OpenStudy (zarkon):

8*(-3)+4*5

OpenStudy (anonymous):

-4?

hero (hero):

Zarkon, it would have been helpful to post a general formula for finding the dot product of two points.

OpenStudy (anonymous):

-100

OpenStudy (zarkon):

\[\vec{a}\cdot \vec{b}=\sum_{i=1}^{n}\left[a_i\times b_i\right]\]

OpenStudy (anonymous):

lost?

OpenStudy (zarkon):

yes the answer is -4

OpenStudy (anonymous):

ok ok

OpenStudy (ash2326):

let A= a1i+a2j+a3k B= b1i+b2j+b3k A.B=a1b1+a2b2+a3b3

OpenStudy (anonymous):

i like the old zarkon avatar